OpenStudy (anonymous):

proof that exp^(x+y) is not equal with (exp^x)(exp^y) given x and y is matrix and xy is not equal to yx please help

OpenStudy (anonymous):

go to this @diran http://en.wikipedia.org/wiki/Matrix_exponential

OpenStudy (anonymous):

may i know k is stand for?

OpenStudy (fwizbang):

Construct a counter example with x = ( 1 0 and y = (0 1 0 -1) 1 0). Since x^2 =y^2 =1, and (x+y)^2 = 2, you can evaluate all the exponential explicitly via Taylor series expansions.